Two cyclists, 79 mi apart, start riding towards each other at the same time. One cycles twice as fast as the other. If they meet 1 hr later, at what average speed is each cyclist traveling?
_____mi/hr(slower cyclist)
_____mi/hr(faster cyclist)

Answer by checkley71(8403)

X+2X=79
3X=79
X=79/3
X=26.33 MPH FOR THE SLOWEST BIKER.
2*26.33=52.67 MPH FOR THE FASTER BIKER.
PROOF
26.33+52.67=79
79=79